OpenStudy (anonymous):

\[\int\frac{dx}{\sqrt{x^2-81}}\]Tried \(\cos u,\sin u,\tan u\) substitutions but nothing fit. :|

OpenStudy (anonymous):

Woops, \(\sec u\) substitution, my bad.

OpenStudy (anonymous):

But what do I do about\[\int\frac{1}{\tan u}\]

OpenStudy (anonymous):

i don't think this is doable with trig-substitution. i don't know. :( sorry.

OpenStudy (anonymous):

It's the first problem in my textbook...

OpenStudy (anonymous):

what calc is this? maybe find it numerically?

OpenStudy (anonymous):

Calc2, non-numerical

OpenStudy (anonymous):

maybe @directrix , @.Sam. , @Callisto, can help?

OpenStudy (callisto):

use sec theta

OpenStudy (anonymous):

Yeah, I just figured it out.

sam (.sam.):

\[x=\sec \theta\]

OpenStudy (anonymous):

Can I ask a further question?

OpenStudy (anonymous):

Or shoudl I create a new question?

sam (.sam.):

yeah, create new one

OpenStudy (callisto):

x=9sec theta ... not sec theta

OpenStudy (anonymous):

I know, it's the general hint that counts.

OpenStudy (anonymous):

New question. http://openstudy.com/study#/updates/4f702f3ce4b0eb858772ee32

sam (.sam.):

\[\sqrt{x^{2}-a^{2}}~~,x=asec \theta\]

OpenStudy (anonymous):

Since I solved it, I think it's safe to say the hint worked. :)